OpenStudy (anonymous):

determine area of triangle. A=112 degrees b=7, and c=11

OpenStudy (dumbcow):

 ok use Law of cosines to find side "a" then Law of Sines to find angle "C" then find height of triangle --> sinC = height/7 then find Area --> Area = 1/2*height*"a"

OpenStudy (dumbcow):

there may be a faster method or formula im not remembering

OpenStudy (anonymous):

how do you find the angles though?

OpenStudy (dumbcow):

\[a^{2} = 7^{2}+11^{2}-2(7)(11)\cos 112\] and \[\frac{\sin C}{11} = \frac{\sin 112}{a}\]

OpenStudy (anonymous):

 CD is the height and is 7 times the sin 68. Then : A = (1/2)bh where AB is the base.

OpenStudy (dumbcow):

see a faster method

OpenStudy (anonymous):

thanks, guys!

OpenStudy (anonymous):

uw!

OpenStudy (mertsj):

Just use the formula : \[A=\frac{1}{2}ab \sin C=\frac{1}{2}(7)(11)\sin 112\]

OpenStudy (mertsj):

A=35.7

OpenStudy (anonymous):

yes, Mertsj's way will also give the same answer because sin 112 = sin 68

OpenStudy (dumbcow):

@BeautyQueen327 , you now have 2 methods, a formula and how that formula is derived i think you should now thoroughly understand how to do this problem :)

OpenStudy (anonymous):

i do! Dont worry! :P I said earlier, thanks, guys! But no one saw it i guess...